# Origins of the fastest stars from merger-disruption of He-CO white dwarfs

### v1

@@ 1.1 HeCO White Dwarfs
More specifically, we investigate the fate of a double degenerate binary system consisting of a primary Helium-rich HeCO WD, with the total mass of made of a CO core of and a massive helium shell of (see details in [24]), and a secondary hybrid HeCO WD with total mass of made of a CO core of and a massive helium shell of (following [34]). Note, that the much higher He shell for this hybrid is attained through a somewhat different evolution than that described in [35], as explored in [24].
@@ 1.2 Hydrodynamical Simulation
We model the merger of two HeCO white dwarfs using the moving-mesh hydrodynamics code Arepo [37, 38, 39, 40]. This code provides a self-gravitating, adaptive mesh framework for simulating astrophysical phenomena, incorporating a Helmholtz equation of state for degenerate matter and a 55 isotope nuclear reaction network to capture the thermonuclear processes involved in the merger. A detailed description of WD merger simulations in Arepo can be found in [24], and our models closely follow similar approaches and modeling. We run our simulation with a cell mass resolution of g, a minimum cell volume of cm, cells, in a total box size of cm to incorporate the explosion and the expansion of the remnant.
@@ 1.2.1 Initial Conditions
Primary WD: Mass = 0.69 M⊙, CO core = 0.59 M⊙, He shell = 0.1 M⊙ as modelled in [24] using MESA [28, 41, 42], and mapped to 3D. Secondary WD: Mass = 0.62 M⊙, CO core = 0.58 M⊙, He shell = 0.038, M⊙, where the composition is based on MESA models in [33]. This model was generated synthetically in the same method as in [36].
@@ 1.2.2 Inspiral and Merger
We then follow the fully realistic evolution of the merger. As the secondary WD approaches the primary, it is tidally deformed and disrupted (s after the end of the accelerated inspiral stage). The dynamical mass transfer of helium from the secondary triggers a helium detonation on the surface of the primary. This detonation propagates around the primary, leading to a secondary detonation in the carbon-oxygen core and a supernova explosion. The partially disrupted secondary WD, still containing its core, is ejected at high velocity by the force of the explosion. These stages are presented in Figure 4 in the SI, while some are also in Figure 1 of the main part. A full movie of the simulation can be found in the SI.
@@ 1.3 Evolution of the Bound Remnant HVWD
To model the long-term evolution of the runaway WD we follow a similar approach as in [7]. We map the output of the simulation ( s after the explosion) to the 1D stellar evolution software MESA. This procedure has been applied before for Athena++ simulations of helium subdwarfs[20] and more recently of helium white dwarfs[8], as well as Arepo simulations of massive white dwarfs [7]. Here, we follow the procedure outlined in reference [7]. The 3D profiles of the runaway WD (taken at the last snapshot, which is 10-20 times dynamical timescales after the explosion, allowing sufficient time for the WD to adjust) are averaged to give a 1D profile with temperature, density, and mass fraction information. A 22 isotope network approx_21_plusco56 is used to model the nuclear reactions during this evolution and includes the isotopes 1H, 3,4He, 12C, 14N, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, 44Ti, 48,56Cr, 52,54,56Fe, 56Co, and 56Ni. It therefore includes the most important reactions of nickel decay and capture on carbon. We relax a white dwarf to our output profiles of composition and entropy, as outlined in [7], and we get rid of almost all of the outer layers where 56Ni dominates the atmosphere to such an extent that the material will become unbound on short timescales (but longer than the hydrodynamical simulation). This cut is chosen again to be the mass coordinate where the energy due to 56Ni decay is twice the gravitational binding energy of the star. This results in a final mass of 0.482 M⊙, with only 0.005 M⊙ cut-off. Once relaxed the white dwarf is allowed to evolve for a maximum time of 100 Myr.
2 Supplementary Information
This supplementary information provides additional details and supporting material for the main text.
2.1 The properties of the hypervelocity WDs
Table 3 summarizes the physical and kinematic properties of the seven fastest hypervelocity WDs, from tables 2 and 3 in ref. [6] (kinematics and radii) and from Refs. [4, 43, 29, 44] for the temperatures and radii.
Name | ||||||||
[km s-1] | [km s-1] | [km s-1] | [K] | [] | [kpc] | [kpc] | [Myr] | |
D6-1 | 120040 | 2045 | 2254 | 7000 | 0.23 | 1.91 | -0.57 | 0.63 |
D6-2 | 8010 | 1151 | 1051 | 7000 | 0.15 | 0.84 | -0.27 | -0.72 |
D6-3 | -2080 | 2248 | 2393 | 7000 | 0.17 | 2.26 | 0.93 | 2.24 |
J1235-3752 | -169410 | 2670 | 2471 | 21000 | 0.104 | 4.53 | 1.73 | 1.87 |
J0927-6335 | -228520 | 2753 | 2519 | 60000 | 0.052 | 4.07 | -0.70 | -1.32 |
J0546+0836 | 120020 | 1699 | 1864 | 95000 | 0.051 | 4.02 | -0.69 | 0.61 |
J1332-3541 | 109050 | 1464 | 1619 | 70000 | 0.017 | 1.63 | 0.74 | 0.64 |
2.2 Detailed evolution of the merger: Movie and snapshots
A movie visualizing the merger of the two HeCO white dwarfs and the subsequent ejection of the HVWD is available online 111https://tinyurl.com/dd2hyperv. The movie shows the tidal disruption of the secondary WD, the double-detonation explosion of the primary, and the high-velocity ejection of the remnant core.
In addition, we provide more detailed snapshots of the evolution in Fig. 4
2.3 Analytic Estimate of Ejection Velocities
The ejection velocity of the HVWD can be estimated analytically by considering the Keplerian velocity of the binary system at the time of the explosion. Assuming that the explosion occurs near the tidal radius, the ejection velocity of the secondary is given by:
where is the gravitational constant, is the total mass of the binary system, and is the tidal radius. The Tidal radius can be approximated as:
Where is the radius of the primary WD and and are the masses of the primary and secondary WDs, respectively. The radii of the WDs are calculated using the mass-radius relation formulation by ref. [45].
Figure 5 shows the expected ejection velocities for a range of binary WD progenitors, calculated using such an analytic estimate. The measured ejection velocity in our simulation matches the analytic prediction well. Nevertheless, this simplified estimate should be taken with a grain of salt, given that the final stages of the tidal evolution and deformation before the SN explosion affect the orbit and the structure of the WDs, and even some of the companion mass is stripped; hence these should only be considered as order of magnitude expectations as to bracket the range of possible ejection velocities in the double HeCO WD merger scenario.
2.4 The properties of HVWD J1332
Given that in both the D6 scenario and our model, the outer layers of the ejected HVWD are stripped, it is unlikely that such a WD would have a hydrogen atmosphere. Therefore, we do not adapt the properties of the HVWD inferred by [29] for J1332 using a hydrogen-rich atmosphere, but the ones assuming a Helium outer atmosphere (hotter and puffier). Nevertheless, for completeness, in Figure 6 we show the HR diagram, radius, and temperature models, and their comparison with observations of J1332 if the hydrogen-rich model of [29] is taken at face value. In that case, the HR position is only consistent within and the inferred radius is no longer consistent with our simulation results. One should note that in that case, J1332 could be consistent with a model of a far more massive CO HVWD, as discussed in [7].
2.5 Population Synthesis Considerations
Population synthesis studies suggest that mergers involving hybrid WDs are not rare among double WD mergers, with as much as 25% of WD mergers including at least one hybrid WD [30, 31]. The mergers of double HeCO WDs represent only a fraction of these, with a smaller fraction likely producing the specific conditions required for the partial disruption and double-detonation explosion observed in our simulation. If such explosions are a significant source of HVWDs, their relatively low occurrence rate could explain the paucity of HVWDs compared to the rate of normal Type Ia SNe.
2.6 Observational Signatures
The peculiar Type Ia SNe produced by the HeCO WD merger scenario are expected to have fainter luminosities and different observational signatures compared to standard SNe Ia. These differences arise from the lower mass of the exploding WD and the distinct composition of the ejecta. Further studies are needed to characterize the detailed observational properties of these explosions and their potential contribution to the observed diversity of SNe Ia. These will be explored in future studies.
3 Software and third party data repository citations

### v2

@@ 1.1 HeCO White Dwarfs
More specifically, we investigate the fate of a double degenerate binary system consisting of a primary Helium-rich HeCO WD, with the total mass of made of a CO core of and a massive Helium shell of (see details in [24]), and a secondary hybrid HeCO WD with total mass of made of a CO core of and a massive Helium shell of (following [40]). See Sec. 2.5 for a discussion on the rate of such interactions. Note that the much higher He shell for this hybrid is attained through a somewhat different evolution than that described in [29], as explored in [24].
@@ 1.2 Hydrodynamical Simulation
We model the merger of two HeCO white dwarfs using the moving-mesh hydrodynamics code Arepo [30, 41, 42, 43]. This code provides a self-gravitating, adaptive mesh framework for simulating astrophysical phenomena, incorporating a Helmholtz equation of state for degenerate matter and a 55-isotopes nuclear reaction network to capture the thermonuclear processes involved in the merger. A detailed description of WD merger simulations in Arepo can be found in [24], and our models closely follow similar approaches and modeling. We run our simulation with a cell mass resolution of g, a minimum cell volume of cm3, cells, in a total box size of cm to incorporate the explosion and the expansion of the remnant.
@@ 1.2.1 Initial Conditions
Primary WD: Mass = 0.69 M⊙, CO core = 0.59 M⊙ (0.26 M⊙ 12C and 0.33 16O), He shell = 0.1 M⊙ as modeled in [24] using MESA [36, 44, 45], and mapped to 3D. Secondary WD: Mass = 0.62 M⊙, CO core = 0.58 M⊙ (equal amount of 12C and 16O), He shell = 0.038, M⊙, where the composition is based on MESA models in [39]. This model was generated synthetically in the same method as in [25].
@@ 1.2.2 Inspiral and Merger
We then follow the fully realistic evolution of the merger. As the secondary WD approaches the primary, it is tidally deformed and disrupted (s after the end of the accelerated inspiral stage). The dynamical mass transfer of Helium from the secondary triggers a Helium detonation on the surface of the primary. This detonation propagates around the primary, leading to a secondary detonation in the Carbon-Oxygen core and a supernova explosion. The partially disrupted secondary WD, still containing its core, is ejected at high velocity by the force of the explosion. These stages are presented in Supplementary Figure 1, while some are also in Figure 1 of the main part. A full movie of the simulation can be found in the SI (Supplementary Video 1).
@@ 1.3 Evolution of the Bound Remnant HVWD
To model the long-term evolution of the runaway WD we follow a similar approach as in [11]. We map the output of the simulation ( s after the explosion) to the 1D stellar evolution software MESA. This procedure has been applied before for Athena++ simulations of Helium subdwarfs[21] and more recently of Helium white dwarfs[12], as well as Arepo simulations of massive white dwarfs [11]. Here, we follow the procedure outlined in reference [11]. The 3D profiles of the runaway WD (taken at the last snapshot, which is 10-20 times dynamical timescales after the explosion, allowing sufficient time for the WD to adjust) are averaged to give a 1D profile with temperature, density, and mass fraction information. A 22 isotope network approx_21_plusco56 is used to model the nuclear reactions during this evolution and includes the isotopes 1H, 3,4He, 12C, 14N, 16O, 20Ne, 24Mg, 28Si, 32S, 36Ar, 40Ca, 44Ti, 48,56Cr, 52,54,56Fe, 56Co, and 56Ni. It therefore includes the most important reactions of nickel decay and capture on Carbon. We relax a white dwarf to our output profiles of composition and entropy, as outlined in [11], and we get rid of almost all of the outer layers where 56Ni dominates the atmosphere to such an extent that the material will become unbound on short timescales (but longer than the hydrodynamical simulation). This cut is chosen again to be the mass coordinate where the energy due to 56Ni decay is twice the gravitational binding energy of the star. This results in a final mass of 0.482 M⊙, with only 0.005 M⊙ cut-off. Once relaxed the white dwarf is allowed to evolve for a maximum time of 100 Myr. We present the results in Figure 3 and Supplementary Figure 2. In addition, as we show in Supplementary Table 2, there is still bound He at the end of our simulation, which is kept there also when following its longer evolution; it might however, keep changing on an even longer timescale. Nevertheless, the existence of a bound He in the remnant will likely lead to He detection in the remnant’s atmosphere.
Data Availability
The MESA input files and inlists for relaxation, heating, and evolution can be accessed at Zenodo [46].
All data underlying this research, which is not in Zenodo (in particular, the many too large, hydrodynamical simulation snapshots), will be available upon reasonable request to the corresponding authors.
Code Availability
@@ Acknowledgments
Corresponding authors: Correspondence and requests for materials should be addressed to Hila Glanz or Hagai Perets.
Competing interests statement
The authors declare no competing interests.